# A kinetic-based regularization method for data science applications

### v1

@@ I Introduction
A common approach for addressing some of these issues is regularization, which can help mitigate overfitting and enhance generalization. Regularization techniques can be divided into two broad categories: non-physics-based and physics-based approaches. Non-physics-based regularization techniques, such as L2 (ridge regression) Hoerl and Kennard (1970) and L1 (lasso) Tibshirani (1996) regularization, have long been standard tools in data science, and aim at reducing the complexity of the model by adding penalties to the model parameters. These techniques involve tuning of hyperparameters through cross-validation or trial and error, which can be costly and time-consuming.
@@
PIML methods help bridging the gap between data-driven learning and the underlying physical principles, enhancing the model’s robustness and ability to generalize beyond the training set. For example, enforcing constraints on preservation of symmetry or conservation laws, allows the model to remain physically consistent, even when making predictions for scenarios not included in the training set. However, hard-wiring physical constraints into ML algorithms can be a challenging task Beucler et al. (2021). Often, soft constraints are employed in the form of penalty terms within the loss function to guide optimization. While this approach is relatively simple, the downside is that it is not guaranteed that the constraints will be strictly enforced in all cases.
@@ II Theory and Methodology
$P(||\tilde{\bm{x}}-\bm{x}_{i}||,\theta)=\frac{p(||\tilde{\bm{x}}-\bm{x}_{i}||,\theta)}{\sum_{j}p(||\tilde{\bm{x}}-\bm{x}_{j}||,\theta)}.$
@@ II.1.3 Second moment
We shall stress that the cancellation of the spurious quadratic term provided by the above expression is exact only when considering points belonging to the training set; nevertheless we can still expect a partial improvement for predictions on previously unseen points.
@@ II.2 Tuning the temperature parameter
In the previous section we did not address how $\theta$, the only free parameter of our method, should be selected. In analogy with statistical physics, where temperature governs the balance between local and global fluctuations, the parameter $\theta$ controls the interaction range between points. A meaningful choice of $\theta$ should scale with a characteristic length $d_{\rm typ}$, which represents the typical spacing between data points. In the low-temperature limit ($\theta\rightarrow 0$), the Gaussian kernel reduces to a Dirac delta function. As a result, each point is effectively isolated, and interactions with neighboring points vanish. This leads to a “cold death” scenario, where the model memorizes individual points (over-fitting) with no meaningful interpolation/generalization for points outside the input dataset. Conversely, in the high-temperature regime ($\theta\gg d_{\rm typ}$), the Gaussian kernel flattens, making all points exert nearly uniform influence. This results in a “hot death”, where the model collapses into a global function, losing its ability to capture fine-grained structures, thus leading to under-fitting. To avoid these extremes, a natural scaling is $\theta=k\cdot d_{\rm typ}$ with $k$ a free scaling parameter, while $d_{\rm typ}$ can be estimated as the median or mean pairwise distance, or the average k-nearest neighbor distance.
@@
This guess can be further refined by employing standard search algorithms Nocedal and Wright (1999).
We propose a simple approach, inspired by the maximum entropy approach to parameter estimation Dempster et al. (1977), in which, starting from Eq. 16, the temperature parameter can be updated iteratively to satisfy the continuum relation:
$\theta^{(n+1)}=\frac{1}{N}\sum_{i=1}^{N}\left(x_{i}-\bar{x}\right)^{2}P(|{\boldsymbol{\xi}}_{i}|,\theta^{(n)}).$ (17)
@@
$\theta^{(n+1)}=\alpha\left[\frac{1}{N}\sum_{i=1}^{N}\left(x_{i}-\bar{x}\right)^{2}P(|{\boldsymbol{\xi}}_{i}|,\theta^{(n)})\right]+(1-\alpha)\theta^{(n)}.$ (18)
@@
The iterative search concludes when the validation error converges, or when the update step falls below a predefined tolerance, or the maximum number of iterations is reached. In this latter case, the value of $\theta$ corresponding to the minimum observed validation error is selected as the optimal parameter.
@@ III Numerical Results
Figure 3: The figure illustrates the progressive improvements in the prediction of the test function, driven by a decreasing energy minimum in the error landscape as the moment correction order increases. Left: RMSE vs. $\theta$ for the 2D Franke function on the validation set $(N_{v})$, using $2000$ randomly sampled points $(N=2000)$ for model training. Right: convergence test for the 2D Franke function on a random test set of $10000$ points, showing the test RMSE as a function of the sampling size $(N)$.
In this section, we evaluate the impact of the moments-based regularization procedure introduced in the previous section for both interpolation and regression tasks. For each test case, we consider a dataset consisting of $N$ points for training the model. This set is then split into a training dataset of size $N_{t}$ and a validation set of size $N_{v}$, such that $N=N_{v}+N_{t}$. These subsets are used to apply the regularization procedure and to determine the optimal value of the parameter $\theta$ (see Sec. II.2). Before training, the data is normalized so that the components of $\bm{x}$ lie within the range $[0,1]$, while the base functional values are scaled to $[-1,1]$. This normalization provides a consistent reference framework and helps comparison of the error committed across different test cases. In what follows we will refer to the following error metrics:
$<L_{1}>=\,<|\hat{\phi}(\bm{x})-\phi(\bm{x})|>,$ (19)
$\rm{RMSE}=\,\frac{1}{\max\,|\phi(\bm{x})|}\,\left(<\left(\hat{\phi}(\bm{x})-\phi(\bm{x})\right)^{2}>\right)^{\frac{1}{2}},$ (20)
$L_{\infty}\,=\,\max\,|\hat{\phi}(\bm{x})-\phi(\bm{x})|.$ (21)
@@ III.1 Effect of moment corrections
We consider a set of $N=2000$ points, randomly picked in the interval $[0,1]\times[0,1]$, subsequently split in a training and test set using a $80:20$ ratio. We apply Algorithm 1, and use the test set for establishing the optimal value of the $\theta$ parameter, or, in other terms, the temperature corresponding to the local minima in the energy landscape of the predicted system.
@@
Turning on a more quantitative ground, in Fig. 3 we investigate the impact of the different correction terms on the overall accuracy, measured in terms of RMSE. In the left panel of the figure, we take as an example the case $N=2000$, and show the energy landscape as a function of the temperature parameter $\theta$. It is instructive to observe how such landscape gets modified when applying constraints to the moments of the distribution. In comparison to the baseline method (blue curve), applying constraints up to the first order of the distribution allows to identify an optimal value of $\theta$ for which we can reduce the error by a factor $\approx 3$, with a further slight improvement when constraining also the second order moment. Ultimately, the constraints carve out a nice local minimum in the energy landscape, which is the reason why they improve the results. In the Figure the right-most values on the x-axis are obtained from the ansatz in Eq. 16. Following the discussion in Sec. II.2, we also show a rescaling of the x-axis, given by $\theta/d_{\rm typ}$, where $d_{\rm typ}$ was taken as the average k-nearest neighbor distance with $k=5$ on the training dataset. This rescaling highlights the regions corresponding to under-fitting and over-fitting, also providing a superior initial guess for identifying the optimal value of $\theta$.
@@
$f_{\text{camel}}(x_{i})=\frac{1}{2(k\sqrt{\pi})^{D}}\left(\exp\left(-\frac{\left(x_{i}-\frac{1}{3}\right)^{2}}{k^{2}}\right)\right.\\ \left.+\exp\left(-\frac{\left(x_{i}-\frac{2}{3}\right)^{2}}{k^{2}}\right)\right),$ (23)
with $k=0.2$.
Table. 1 presents the results of our analysis, comparing the accuracy of the baseline model against the regularized one across different dimensions and dataset sizes. We observe a consistent improvement in accuracy when moment corrections are applied. The extent of this improvement (fifth column in the table), strongly depends on both $D$ and $N$: bigger gains can be observed for increasingly large datasets, but the trend is slowed down when also the dimensionality $D$ is increased, as consequence of the so called curse of dimensionality.
@@
$D$ $N$ Without Correction With Correction Ratio
@@
Table 1: Convergence of RMSE for $f_{\text{camel}}(\bm{x})$ across different dimensions ($D$) and training sizes ($N$) of random points. The RMSE is computed using 10000 randomly sampled points ($\bm{x}\in[0,1]^{D}$) outside the training set. The ratio column represents the improvement due to the correction.
@@ III.2 Interpolation of a fractal curve
We observe that uniform sampling benefits the RBF interpolation, which consistently outperforms the present method to varying degrees, particularly in cases with lower complexity and sparser sampling. As both the complexity and the number of sampling points increase, the performance gap narrows, with the present method progressively catching up to the RBF interpolation. However, the trend reverses in the case of random sampling. The plain RBF struggles in comparison to the present method, as can be observed in the bottom panel of Fig. 4. The reason come from the fact that when data is sampled randomly, clusters of data points naturally arise, resulting in an uneven distribution across the domain. Such clustering leads to ill-conditioned interpolation matrix, since certain regions become oversampled while others remain under-sampled, in turn leading to numerical instabilities and inaccuracy.
@@
Figure 5: Comparison of 2D Franke function predictions for $5\%$ Gaussian noise ($s=0.05$) trained and tested on random sets of $1,000$ and $10,000$ points, respectively. Top panel: (noisy) training data. Middle panel: RBF prediction. Bottom panel: prediction with present method.
@@ III.3 Regression on noisy data
where $\epsilon$ is a random variable drawn from a Gaussian distribution, and $s$ is a scale factor that controls the noise intensity. For example, $s=0.05$ corresponds to $5\%$ noise in the data. The Gaussian noise is added with $\sigma=1/3$ and $\mu=0$, based on the $6\,\sigma$ rule. This ensures that most data points fall within the range $[-1,1]$ ($\mu\pm 3\sigma$), with only a small number of outliers. Since the focus is on de-noising the data, the goal is to minimize the error relative to the (generally unknown) base function on the test set. This represents the hidden minima, which remains unknown to the model, while the training is carried out using the noisy data itself.
@@
The effect of the regularization is shown more clearly in Fig. 6, where we investigate the impact of the correction terms on the overall accuracy. The plot in the left panel shows the energy landscape as a function of the temperature parameter $\theta$ for $s=0.05$. The continuous lines show the error on the test data, while the dotted lines show the error with respect to the base (noise-less) data. When moment-based corrections are applied, we observe that the error on the noisy test data not only decreases compared to the case without corrections, but also allows to access a better local minima in the error-curve for the noise-less data. As the signal to noise ration is decreased, the advantadge of the moments correction is weakened, as shown in the right panel of Fig. 6, where the same analysis is repeated for $s=0.2$; say that at least it does not do worse than the baseline method without corrections.
@@
Figure 8: The top three panels show the ground truth, i.e., the noiseless numerical data which we aim to reproduce and the results obtained using Delaunay triangulation (Linear) and the present method respectively; here $25\,\%$ of data points were used for the training. The bottom panels show the $<L_{1}>$, $RMSE$, and $L_{\infty}$ error norms, while increasing the amount of data used to train the model. We compare the present method with linear and cubic interpolation methods with a noise level of $(s=0.15)$. The error metrics are computed using the data not supplied in the training set.
@@
As a final example, we consider a dataset derived from numerical simulations of convective fluid flows, where the underlying analytical form is unknown. This case study illustrates a typical example of application, where a smaller, finite set of data points is used to generate data at a finer scale. A similar situation arises when dealing with experimental data, which is often sparse and noisy. In such cases, interpolation methods need to be both fast and robust, particularly when the aim is to generate more accurate, noise-free data from a limited dataset.
For this study, we use a $240\times 240$ dataset from a snapshot of a 2D temperature field obtained from numerical simulations. For model training, $25\%$ of the dataset is randomly selected and split into training and validation sets in an $80:20$ ratio. In other words, one-fourth of the data is used for training, while the remaining three-fourths are reserved for testing the model’s predictions. To mimic a noisy dataset, we use Eq. 25 to introduce varying degrees of Gaussian noise. We compare our method with two commonly used interpolation techniques: linear and cubic interpolation. For linear interpolation, we use Delaunay triangulation Chen and chao Xu (2004), where the value at an unknown point is estimated by performing piecewise linear interpolation based on the nearest known points that form a triangle. For cubic interpolation, we employ the Clough-Tocher method Clough and Tocher (1965), which fits cubic polynomials ($C^{1}\text{ continuous}$) to the triangles, resulting in a smoother interpolation compared to the linear method, which only ensures $C^{0}\text{ continuity}$ (i.e., piecewise linear).
@@
Considering now the case with $s=0.15$ (for which the top panel in Fig. 8 shows a qualitative comparison), we evaluate the effect of increasing the number of points forming the training set. The bottom panel show the trend for the three different error metrics $<L_{1}>$, $RMSE$, and $L_{\infty}$, while increasing the fraction of training points. The present method outperforms the interpolator schemes on all the metrics taken in exam. Moreover it is interesting to observe from the analysis of the RMSE that linear and cubic reach already plateu when $10\%$ of the data is used, adding more does not help. On the other hand, adding more data with the present method helps improving well beyond that point.
Considering now the case with $s=0.15$, for which top panel in Fig. 8) offers a visual representation, we evaluate the effect of increasing the number of points in the training set. The bottom panel shows the trend for three different error metrics: $<L_{1}>$, RMSE, and $L_{\infty}$, as the fraction of training points increases. The present method outperforms the interpolation schemes across all metrics. Notably, the RMSE analysis reveals that both the linear and cubic interpolation methods reach a plateau already when only $10\%$ of the data is used, meaning adding more data does not improve their performance. In contrast, the present method continues to improve well beyond that point, allowing a more effective use of additional data.

### v2

@@ I Introduction
A common approach for addressing some of these issues is regularization, which can help mitigate overfitting and enhance generalization. Regularization techniques can be divided into two broad categories: non-physics-based and physics-based approaches. Non-physics-based regularization techniques, such as L2 (ridge regression) Hoerl and Kennard (1970) and L1 (lasso) Tibshirani (1996) regularization, have long been standard tools in data science, and aim at reducing the complexity of the model by adding penalties to the model parameters. These techniques involve tuning of hyperparameters through cross-validation or trial and error, which can be costly and time-consuming. A more efficient and systematic approach is Maximum Likelihood Estimation (MLE), which chooses hyperparameters by maximizing the marginal likelihood Rasmussen and Williams (2006).
@@
PIML methods help bridging the gap between data-driven learning and the underlying physical principles, enhancing the model’s robustness and ability to generalize beyond the training set. For example, enforcing constraints on preservation of symmetry or conservation laws, allows the model to remain physically consistent, even when making predictions for scenarios not included in the training set. Often, the kernels used in these models are also inspired by physics. In Higher Dimensional Model Representation using Gaussian Process Regression (HDMR-GPR) Rabitz and Alis (1999), for instance, the kernel is constructed as a sum of individual Gaussian functions corresponding to increasing orders of interaction, an approach inspired by many-body interactions in quantum and statistical mechanics. However, hard-wiring physical constraints into ML algorithms can be a challenging task Beucler et al. (2021). Often, soft constraints are employed in the form of penalty terms within the loss function to guide optimization. While this approach is relatively simple, the downside is that it is not guaranteed that the constraints will be strictly enforced in all cases.
@@ II Theory and Methodology
[…]
@@ II.1.3 Second moment
It is important to note that the cancellation of the spurious quadratic term in the above expression is exact only for points in the training set (and provided that $\Theta$ has no spatial dependence). Nevertheless, a partial improvement can still be expected for predictions on unseen points, as we will show in the coming sections (see also Appendix B for more details).
@@ II.2 Tuning the temperature parameter
In the previous section we did not address how $\theta$, the only free parameter of our method, should be selected. In analogy with statistical physics, where temperature governs the balance between local and global fluctuations, the parameter $\theta$ controls the interaction range between points. A meaningful choice of $\theta$ should scale with the square of a characteristic length $d_{\rm typ}$, which represents the typical spacing between data points. In the low-temperature limit ($\theta\rightarrow 0$), the Gaussian kernel reduces to a Dirac delta function. As a result, each point is effectively isolated, and interactions with neighboring points vanish. This leads to a “cold death” scenario, where the model memorizes individual points (over-fitting) with no meaningful interpolation/generalization for points outside the input dataset. Conversely, in the high-temperature regime ($\theta\gg d^{2}_{\rm typ}$), the Gaussian kernel flattens, making all points exert nearly uniform influence. This results in a “hot death”, where the model collapses into a global function, losing its ability to capture fine-grained structures, thus leading to under-fitting. To avoid these extremes, a natural scaling is $\theta=k\cdot d^{2}_{\rm typ}$ with $k$ a free scaling parameter, while $d_{\rm typ}$ can be estimated as the median or mean pairwise distance, or the average k-nearest neighbor distance.
@@
Starting from this guess one can determine an optimal value for $\theta$ employing standard search algorithms Nocedal and Wright (1999).
We propose here a simpler approach, which does not involve computation of gradients, inspired by the maximum entropy approach to parameter estimation Dempster et al. (1977), in which, starting from Eq. 16, the temperature parameter can be updated iteratively to satisfy the continuum relation:
[…]
@@
[…]
@@
The iterative search concludes when the validation error converges, when the update step falls below a predefined tolerance, or the maximum number of iterations is reached. In this latter case, the value of $\theta$ corresponding to the minimum observed validation error is selected as the optimal parameter. In Appendix D we compare this approach with the Maximum Likelihood Estimate approach.
@@ III Numerical Results
Figure 3: The figure illustrates the progressive improvements in the prediction of the test function, driven by a decreasing energy minimum in the error landscape as the moment correction order increases. Left: RMSE vs. $\theta$ for the 2D Franke function on the validation set $(N_{v})$, using $2000$ randomly sampled points $(N=2000)$ for model training. Right: convergence test for the 2D Franke function on a test set of $10000$ uniformly spread points, showing the test RMSE as a function of the sampling size $(N)$.
In this section, we evaluate the impact of the moments-based regularization procedure introduced in the previous section for both interpolation and regression tasks. For each test case, we consider a dataset consisting of NN points for training the model. This set is then split into a training dataset of size $N_{t}$ and a validation set of size $N_{v}$, such that $N=N_{v}+N_{t}$. These subsets are used to apply the regularization procedure and to determine the optimal value of the parameter $\theta$ (see Sec. II.2). Before training, the data is normalized so that the components of $\bm{x}$ lie within the range $[0,1]$, while the base functional values are scaled to [1,1][-1,1]. This normalization provides a consistent reference framework and helps comparison of the error committed across different test cases. Alternatively, one can consider instead a dimensionless representation, scaling quantities by appropriate power of $d_{\rm typ}$ (e.g. $\bm{\xi}_{i}/d_{\rm typ}$, $\theta/d^{2}_{\rm typ}$, etc).
In what follows we will refer to the following error metrics ($\hat{\phi}(\bm{x})$ being the predicted value):
$\langle L_{1}\rangle=\;\bigl{\langle}|\hat{\phi}(\bm{x})-\phi(\bm{x})|\bigr{\rangle},$ (19)
$\mathrm{RMSE}=\;\frac{1}{\max|\phi(\bm{x})|}\,\Bigl{\langle}\bigl{(}\hat{\phi}(\bm{x})-\phi(\bm{x})\bigr{)}^{2}\Bigr{\rangle}^{\tfrac{1}{2}},$ (20)
$L_{\infty}\;=\;\max\,|\hat{\phi}(\bm{x})-\phi(\bm{x})|,$ (21)
We also introduce the coefficient of determination $R^{2}$, defined as:
$R^{2}=1-\frac{\left\langle\bigl{(}\hat{\phi}(\bm{x})-\phi(\bm{x})\bigr{)}^{2}\right\rangle}{\left\langle\bigl{(}\phi(\bm{x})-\langle\phi(\bm{x})\rangle\bigr{)}^{2}\right\rangle}.$ (22)
@@ III.1 Effect of moment corrections
We consider a set of $N=2000$ points, randomly picked in the interval $0,1]\times[0,1]$, subsequently split in a training and test set using a $80:20$ ratio. We apply Algorithm 1, and use the test set for establishing the optimal value of the $\theta$ parameter, or, in other terms, the temperature corresponding to the local minima in the energy landscape of the predicted system.
@@
Turning on a more quantitative ground, in Fig. 3 we investigate the impact of the different correction terms on the overall accuracy, measured in terms of RMSE. In the left panel of the figure, we take as an example the case $N=2000$, and show the energy landscape as a function of the temperature parameter $\theta$. It is instructive to observe how such landscape gets modified when applying constraints to the moments of the distribution. In comparison to the baseline method (blue curve), applying constraints up to the first order of the distribution allows to identify an optimal value of $\theta$ for which we can reduce the error by a factor $\approx 3$, with a further slight improvement when constraining also the second order moment. Ultimately, the constraints carve out a nice local minimum in the energy landscape, which is the reason why they improve the results. In the Figure the right-most values on the x-axis are obtained from the ansatz in Eq. 16. Following the discussion in Sec. II.2, we also show a rescaling of the x-axis, given by $\theta/d^{2}_{\rm typ}$, where $d_{\rm typ}$ was taken as the average k-nearest neighbor distance with $k=5$ on the training dataset. This rescaling highlights the regions corresponding to under-fitting and over-fitting, also providing a superior initial guess for identifying the optimal value of $\theta$.
@@
$f_{\text{camel}}(\mathbf{x})=\frac{1}{2(k\sqrt{\pi})^{D}}\left(\exp\left(-\sum_{i=1}^{D}\frac{\left(x_{i}-\frac{1}{3}\right)^{2}}{k^{2}}\right)\right.\\ \left.+\exp\left(-\sum_{i=1}^{D}\frac{\left(x_{i}-\frac{2}{3}\right)^{2}}{k^{2}}\right)\right),$ (24)
where $\mathbf{x}=(x_{1},x_{2},......,x_{D})\in\mathbb{R}^{D}$, with $k=0.2$. Note that this function features a strong unbounded multiplicative coupling.
Table. 1 presents the results of our analysis, comparing the accuracy of the baseline model against the regularized one across different dimensions and dataset sizes. We observe a consistent improvement in accuracy when moment corrections are applied. The extent of this improvement (reported in the fifth column of the table), strongly depends on both $D$ and $N$: bigger gains can be observed for increasingly large datasets, but the trend slows down with increasing dimensionality $D$, as consequence of the so called curse of dimensionality.
@@
$D$ NN Without Correction With Correction Ratio
@@
Table 1: Convergence of RMSE for $f_{\text{camel}}(\bm{x})$ across different dimensions ($D$) and training sizes (NN) of random points. The RMSE is computed using 10000 randomly sampled points ($\bm{x}\in[0,1]^{D}$) outside the training set. The ratio column represents the improvement due to the correction.
To further investigate the method, we consider another example of a high-dimensional function: the 6D Ackley function, which exhibits weaker coupling between variables compared to the Camel function in the same number of dimensions:
$f_{\text{ackley}}(\mathbf{x})=-20\exp\left(-0.2\sqrt{\frac{1}{6}\sum_{i=1}^{6}x_{i}^{2}}\right)\\ -\exp\left(\frac{1}{6}\sum_{i=1}^{6}\cos(2\pi x_{i})\right)+20+\exp(1),$ (25)
where $\mathbf{x}=(x_{1},x_{2},......,x_{6})\in\mathbb{R}^{6}$. As with the 6D Camel function, we observe that the proposed correction scheme progressively improves the accuracy as the grid becomes denser, as summarized in Table 2. This motivates a closer examination of the method’s performance in the sparse-data regime, where training points are limited. A detailed analysis of the trade-offs between accuracy and computational cost, alongside with comparisons to other methods, is provided in Appendix C.
RMSE
$D$ NN Without Correction With Correction Ratio
6 2000 $2.91\times 10^{-2}$ $2.50\times 10^{-2}$ 1.16
6 4000 $2.47\times 10^{-2}$ $2.01\times 10^{-2}$ 1.22
6 8000 $2.02\times 10^{-2}$ $1.46\times 10^{-2}$ 1.38
6 16000 $75\times 10^{-2} 1.1$$10\times 10^{-2}$ 1.59
6 32000 $1.43\times 10^{-2}$ $8.68\times 10^{-3}$ 1.65
6 64000 $1.13\times 10^{-2}$ $6.27\times 10^{-3}$ 1.80
6 100000 $9.98\times 10^{-3}$ $4.27\times 10^{-3}$ 2.33
Table 2: Convergence of RMSE for $6D$ $f_{\text{ackley}}(\bm{x})$ across different training sizes (NN) of random points. The RMSE is computed using 10000 randomly sampled points ($\bm{x}\in[0,1]^{6}$) outside the training set. The ratio column represents the improvement due to the correction.
@@ III.2 Interpolation of a fractal curve
We observe that uniform sampling benefits the RBF interpolation, which consistently outperforms the present method to varying degrees, particularly in cases with lower complexity and sparser sampling. As both the complexity and the number of sampling points increase, the performance gap narrows, with the present method progressively catching up to the RBF interpolation. However, the trend reverses in the case of random sampling. The plain RBF struggles in comparison to the present method, as can be observed in the bottom panel of Fig. 4. The reason comes from the fact that when data is sampled randomly, clusters of data points naturally arise, resulting in an uneven distribution across the domain. Such clustering leads to ill-conditioned interpolation matrix, since certain regions become oversampled while others remain under-sampled, in turn leading to numerical instabilities and inaccuracy.
@@
Figure 5: Comparison of 2D Franke function predictions for $5\%$ Gaussian noise ($s=0.05$) trained and tested on a random set of $1000$ and a uniform set of $10000$ points respectively. Top panel: (noisy) training data. Middle panel: RBF prediction. Bottom panel: prediction with present method.
@@ III.3 Regression on noisy data
where $\epsilon$ is a random variable drawn from a Gaussian distribution, and $s$ is a scale factor that controls the noise intensity. For example, $s=0.05$ corresponds to $5\%$ noise in the data. The Gaussian noise is added with $\sigma=1/3$ and $\mu=0$, based on the $6\,\sigma$ rule. This ensures that most data points fall within the range [1,1][-1,1] ($\mu\pm 3\sigma$), with only a small number of outliers. Since the focus is on de-noising the data, the goal is to minimize the error relative to the (generally unknown) base function on the test set. This represents the hidden minima, which remains unknown to the model, while the training is carried out using the noisy data itself.
@@
Figure 8: The top three panels show the ground truth, i.e., the noiseless numerical data which we aim to reproduce and the results obtained using Delaunay triangulation (Linear) and the present method respectively; here $25\,\%$ of data points were used for the training. The bottom panels show the $L_{1}>$, $\rm RMSE$, and $L_{\infty}$ error norms, while increasing the amount of data used to train the model. We compare the present method with linear and cubic interpolation methods with a noise level of $(s=0.15)$. The error metrics are computed using the data not supplied in the training set.
Figure 9: The top three panels show the ground truth, i.e., the noiseless recorded temperature anomaly data (from Copernicus Climate Change Service) which we aim to reproduce and the results obtained using Clough-Tocher triangulation (Cubic) and the present method respectively; here $25\,\%$ of data points were used for the training. The bottom panels show the $\rm{RMSE}$ and the coefficient of determination $R^{2}$ with respect to the ground truth, while increasing the amount of data used to train the model. We compare the present method with linear and cubic interpolation methods with a noise level of $(s=0.35)$. The error metrics are computed using the data not supplied in the training set.
@@
Next, we consider a dataset derived from numerical simulations of convective fluid flows, where the underlying analytical form is unknown. This case study illustrates a typical example of application, where a smaller, finite set of data points is used to generate data at a finer scale. A similar situation arises when dealing with experimental data, which is often sparse and noisy. In such cases, interpolation methods need to be both fast and robust, particularly when the aim is to generate more accurate, noise-free data from a limited dataset.
For this study, we use a $240\times 240$ dataset from a snapshot of a 2D temperature field obtained from numerical simulations of phase-change convection (see Ref. Huber et al. (2008) for details on the numerical setup). For model training, $25\%$ of the dataset is randomly selected and split into training and validation sets in an $80:20$ ratio. In other words, one-fourth of the data is used for training, while the remaining three-fourths are reserved for testing the model’s predictions. To mimic a noisy dataset, we use Eq. 27 to introduce varying degrees of Gaussian noise Wei et al. (2023); Sadeghi et al. (2020). We compare our method with two commonly used interpolation techniques: linear and cubic interpolation. For linear interpolation, we use Delaunay triangulation Chen and chao Xu (2004), where the value at an unknown point is estimated by performing piecewise linear interpolation based on the nearest known points that form a triangle. For cubic interpolation, we employ the Clough-Tocher method Clough and Tocher (1965), which fits cubic polynomials ($C^{1}\text{ continuous}$) to the triangles, resulting in a smoother interpolation compared to the linear method, which only ensures $C^{0}\text{ continuity}$ (i.e., piecewise linear).
@@
Considering now the case with $s=0.15$, for which top panel in Fig. 8 offers a visual representation, we evaluate the effect of increasing the number of points in the training set. The bottom panel shows the trend for three different error metrics: $L_{1}>$, RMSE, and $L_{\infty}$, as the fraction of training points increases. The present method outperforms the interpolation schemes across all metrics. Notably, the RMSE analysis reveals that both the linear and cubic interpolation methods reach a plateau already when only $10\%$ of the data is used, meaning adding more data does not improve their performance. In contrast, the present method continues to improve well beyond that point, allowing a more effective use of additional data.
As a final example, more closely aligned with real-world applications involving experimental data, we consider observational data from the Climate Data Store (CDS) of the Copernicus Climate Change Service (C3S) Copernicus Climate Change Service (2024) (C3S), focusing on the European region. The dataset comprises 153 latitude points and 261 longitude points as spatial coordinates. The ground truth, shown in the first figure of the top panel in Fig. 9, represents the near-surface (approximately 2m) temperature anomaly for the year 2016. This anomaly is computed as the difference between the pointwise temperature in 2016 and the long-term average temperature from the 1991-2020 climate reference period. We inject varying degrees of Gaussian noise to the data Zur et al. (2009); Chen et al. (2022). Figure 9 illustrates the case with $35\%$ noise, higher than in the previous example, to explore the method’s robustness under increased uncertainty. As seen in the second and third plots of the top panel, the proposed method yields a visibly smoother and more accurate reconstruction compared to the Clough-Tocher approach. This improvement is further supported by the bottom panel of Fig. 9, which shows lower RMSE and higher $R^{2}$ scores across increasing training fractions, clearly outperforming both Delaunay and Clough-Tocher methods.